Problem Page
Ammonium phosphate is an important ingredient in many fertilizers. It can be made by reacting phosphoric acid with ammonia .
What mass of ammonium phosphate is produced by the reaction of 2.8 gof phosphoric acid?

Round your answer to significant digits.
1
Expert's answer
2019-02-25T08:18:56-0500

2NH3 + H3PO4 = (NH4)2HPO4

M (H3PO4) = 3 + 31 + 64 = 98 g/mol

n (H3PO4) = 2.8/98 = 0.029 mole

n (H3PO4) = n ((NH4)2HPO4)

M ((NH4)2HPO4) = (14+4)*2+1+31+16*4 = 132 g /mol

m ((NH4)2HPO4) = 0.029*132 = 3.8 g